# Denoising and Reconstruction of Nonlinear Dynamics using Truncated Reservoir Computing

### v1

@@ V Discussion
The above examples were focused on showing the performance of RC in discarding additive noise. The denoising performance is examined for the smooth dynamics of the Lorenz attractor and the firing patterns of the AdEx model. While the RC performs well in both cases, it gives better accuracy for smooth dynamics. This issue is rooted in the mathematical formulation of ESN, where the state is expressed as a linear combination of the state at the preceding time step and linear subspaces acting on the state and input instances using a smooth activation function. Perhaps, a different activation function (e.g., Sigmoid, ReLU), which can better approximate discontinuities based on the input instances, may help to build a new RC architecture tailored for non-smooth dynamics. In this respect, building the RC based on the spiking neuro-physical models like a liquid state machine could provide additional flexibility for simulating bursting nonlinear dynamics Hazan and Manevitz (2012). Therefore, the choice of mathematical expression used for constructing the RC appears to be rather problem-specific.
Another important feature of the RC is the generalization behavior when the dynamical parameters change. It is shown that, in a relatively wide neighborhood of the parameters used for training, the denoising gain remains above one. Including extra training data sets with varying dynamical parameters allows for acquiring higher accuracy when dealing with potential parameter changes. However, having a larger training data set creates additional computational costs, and when a hyperparameter optimization is required, the computational cost scales up directly with the cube of the size of the data set due to the matrix inversion in Eq. 4. Therefore, it is important to select the parameters carefully for generating suitable training data and avoid unnecessary computations.
The denoising performance varies based on the SNR. From the Lorenz attractor example, it was seen that the reservoir can decrease or increase the noise in the data. The former often occurs at lower SNRs, where the denoising gain is considerably greater than one. On the other hand, for very high SNRs, the RC cannot distinguish the low-amplitude noise from dominant nonlinear dynamics, and as a result, the RC produces adverse denoising results (see FIG. 3). Therefore, when handling denoising tasks, adequate care should be given to avoid misusing the RC.
The pruning method applied for truncating the reservoir nodes works when the number of reservoir nodes is relatively small (less than 1,000 in this work). The computational efforts required for pruning redundant nodes increase with the number of reservoir nodes. This limitation arises from the use of a simple greedy approach, which explores all nodes to identify redundant ones based on a preselected performance criterion like MSE. However, this limitation could be relaxed by using graph-based optimization techniques, where the reservoir nodes and edges can be reduced globally without having to delve into checking specific points Scarselli et al. (2009).
The frequency content of the noise also affects the denoising performance. The RC provides good results for GWN and violet noise processes and mitigates high-frequency noise components significantly better than low-frequency ones. This conclusion was strengthened when the RC failed to remove additive pink noise from the data, wherein the low-frequency components are more dominant. This negative performance is expected for the low-frequency range since the noise becomes more entangled with the nonlinear dynamics, and therefore, the RC will be unable to distinguish noise from the signal. Thus, the primary hypothesis of this paper needs slight refinement in terms of clarifying the denoising merits, which appear to be limited to discarding high-frequency noise at low SNR values. This issue opens up new directions toward combining the RC with the off-the-shelf wavelet scattering Mallat (2012) and less-studied nonlinear phase-space-based filtering techniques like the GHKSS Grassberger et al. (1993). However, these approaches were not exercised herein, as investigating the potential of RC was the primary focus of this work.
@@ VI Conclusion
A new RC framework was proposed for the reconstruction and denoising of nonlinear dynamics from noisy measured data. The performance of the computing reservoir was investigated in terms of the noise intensity, the sensitivity to dynamical parameters, and the noise frequency content. The optimization of the hyperparameters, the reservoir nodes, the input connectivity, the readout nodes, and the ridge regression coefficient was proven to be necessary and effective to attain a well-trained reservoir computer. Reducing the reservoir dimensionality and removing ineffective edges through a pruning technique significantly improved denoising accuracy, surpassing the performance of the classical reservoir with a fixed number of nodes and random connectivity. At high noise levels, the proposed RC outperformed an EKF implementation. This superior performance was achieved despite the fact that EKF had an advantage as it used precise physical knowledge through the state-space model, while the RC implicitly learned the dynamics solely from the training data.
@@ Acknowledgements.
This research has received funding from the Australian Research Council Linkage Project under grant number LP200301196. Additionally, the first draft of this manuscript was prepared during O. Sedehi’s visit to TU Berlin, supported by Key Technology Partnership (KTP) SEED funding scheme. This financial support is gratefully appreciated.
@@
The data that support the findings of this study are available from the corresponding author upon reasonable request.
@@ Appendix A Extended Kalman Filter
It is assumed that the dynamical system under consideration can be described in discrete-time through a nonlinear state-space relationship given as:
@@
$\mathbf{x}_{k|k-1}=\mathbf{f}(\mathbf{x}_{k-1|k-1},\mathbf{u}_{k})$ (10)
@@ Appendix B Nonlinear equations of Lorenz system
where $\sigma$ is the Prandtl number, $\rho$ is Rayleigh number, and $\beta$ is dissipation rate. Note that selecting different parameters can entirely change the dynamic regimes. The dynamics of the Lorenz system are nonlinear, chaotic, and deterministic, which depend heavily on the initial conditions.
@@ Appendix C Mathematical formulation of AdEx model
where $V$ is voltage, $I$ is input current, $w$ is adaptation parameter, $V_{r}$ is leak reversal potential, $t_{f}$ is firing times at which the voltage exceeds $V_{T}$ threshold, $\delta(.)$ is Dirac delta function. Other parameters are constant coefficients, including $\tau_{m}=5ms$ is membrane time scale, $\tau_{w}=100ms$ is adaptation time constant, $R=500M\Omega$ is membrane resistance, $V_{r}=-55mV$ is rest potential, $V_{T}=-51mV$ is voltage threshold, $\Delta_{T}=-2mV$ is sharpness threshold, $a=-0.5ns$ is adaptation voltage coupling, and $b=7pA$ is spike-triggered adaptation increment. At spike time $t=t_{f}$, there will be a sharp rise in the adaptation parameter equal to $b$, as described via the Dirac delta function.

### v2

@@ V Discussion
The examples above primarily demonstrate the RC system’s ability to discard additive noise. Denoising performance is evaluated using both the smooth dynamics of the Lorenz attractor and the spiking patterns of the AdEx model. While RC performs well in both scenarios, it achieves higher accuracy with smooth dynamics. This is attributed to the mathematical formulation of RCs, where the state is expressed as a linear combination of the previous state and input using a smooth activation function.
To better handle discontinuities, alternative activation functions may be needed, ones that can approximate abrupt changes more efficiently. In this context, architectures based on spiking neuro-physical models, such as liquid state machines, could offer greater flexibility for simulating bursting nonlinear dynamics Hazan and Manevitz (2012). The mathematical structure of RC appears to be problem-specific and should be tailored accordingly.
The denoising performance also varies with the SNR. For lower SNRs, the RC significantly reduces noise, yielding a denoising gain well above unity. However, at very high SNRs, the RC struggles to distinguish low-amplitude noise from dominant nonlinear dynamics, sometimes resulting in adverse denoising effects (see FIG. 3). Therefore, adequate care should be given to avoid misusing the RC.
The frequency characteristics of noise play a significant role in the denoising performance of reservoir computing. The RC demonstrates strong results when dealing with GWN and violet noise, effectively suppressing high-frequency components. However, its performance deteriorates with pink noise, where low-frequency components dominate. This limitation is expected, as low-frequency noise tends to overlap with the underlying nonlinear dynamics, making it difficult for the RC to distinguish signal from noise. This observation suggests that the denoising capabilities of RC are most effective in scenarios involving high-frequency noise at low SNR levels. Consequently, the primary hypothesis of this study requires slight refinement to reflect that RC’s denoising strengths are frequency-dependent.
These findings open up new avenues toward combining the RC with wavelet scattering Mallat (2012) and less-studied nonlinear phase space-based filtering methods like the GHKSS approach Grassberger et al. (1993). While these hybrid strategies were not exercised herein, they represent compelling directions for enhancing RC’s performance in more complex noise environments.
Another important feature of the RC is its generalization capability when the dynamical parameters change. The RC maintains its denoising gain above one across a relatively wide neighborhood of the training parameters. Including extra training data sets with varying dynamical parameters allows for acquiring higher accuracy when dealing with potential parameter changes. This adaptability is crucial for real-world applications where system dynamics may not remain constant.
While expanding the training data set improves generalization, it also increases computational costs. Specifically, when a hyperparameter optimization is required, the cost scales cubically with the dataset size due to matrix inversion in Eq. 4. Therefore, careful selection of training parameters is essential to balance accuracy and computational efficiency.
The computational cost of the ridge regression is proportional to $O(N^{3})$ due to matrix multiplication for calculating the reservoir states, where $N$ is the number of reservoir nodes. The ridge regression requires a matrix multiplication and inversion of the same dimension as the reservoir matrix, having a computational cost proportional to $O(N^{3})$ if the Gauss-Jordan approach is employed for the matrix inversion Golub and Loan (2013). Pruning reservoir nodes can reduce the dimensionality, and if the number of pruned nodes is $M$, the computational complexity can be characterized as $O((N-M)^{3})$. This reduction can be considerably large depending on the number of pruned nodes. However, this assessment of the computational cost requires accounting for the computational efforts related to pruning the reservoir. In fact, at the training stage, the cost of pruning nodes and edges is considerable as it scales up with the number of iterations required to search for $M$ ineffective reservoir nodes.
The pruning method used to reduce the number of reservoir nodes is effective only when the node count is relatively small. The current greedy approach identifies redundant nodes based on a performance criterion like MSE, but its scalability is limited. This limitation could be improved by adopting graph-based optimization techniques, which allow for global reduction of reservoir nodes and edges without exhaustive point-by-point evaluation Scarselli et al. (2009).
@@ VI Conclusion
A new RC framework was proposed for the reconstruction and denoising of nonlinear dynamics from noisy measured data. The performance of the computing reservoir was investigated in terms of the noise intensity, the sensitivity to dynamical parameters, and the noise frequency content. The optimization of the hyperparameters, the reservoir nodes, the input connectivity, the readout nodes, and the ridge regression coefficient was proven to be necessary and effective to attain a well-trained reservoir computer. Reducing the reservoir dimensionality and removing ineffective edges through a pruning technique significantly improved denoising accuracy, surpassing the performance of the classical reservoir with a fixed number of nodes and random connectivity. At high noise levels, the proposed RC outperformed an EKF implementation. This superior performance was achieved despite the fact that the EKF had an advantage as it used precise physical knowledge through the state-space model, while the RC implicitly learned the dynamics solely from the training data. Since the EKF is a near-optimal filter, having observed superior performance for the RC is of great importance from an optimality standpoint.
@@ Acknowledgements.
This research has received funding from the Australian Research Council Linkage Project under grant number LP200301196. Additionally, the first draft of this manuscript was prepared during O. Sedehi’s visit to TU Berlin, supported by the Key Technology Partnership (KTP) SEED funding scheme. This financial support is gratefully appreciated.
@@
All codes and data that support the findings of this study are openly available in GitHub at “https://github.com/Omid-Sedehi/ReservoirComputer.”
@@ Appendix A Extended Kalman filter
It is assumed that the dynamical system under consideration can be described in discrete-time through a nonlinear state-space relationship given as
@@
$\mathbf{x}_{k|k-1}=\mathbf{g}(\mathbf{x}_{k-1|k-1},\mathbf{u}_{k})$ (10)
@@ Appendix B Nonlinear equations of Lorenz system
where $\sigma$ is the Prandtl number, $\rho$ is Rayleigh number, and $\beta$ is dissipation rate. Note that selecting different parameters can entirely change the dynamic regimes. The dynamics of the Lorenz system are nonlinear, chaotic, and smooth deterministic, which depend heavily on the initial conditions.
@@ Appendix C Mathematical formulation of AdEx model
where $V$ is voltage, $I$ is input current, $w$ is adaptation parameter, $V_{r}$ is leak reversal potential, $t_{f}$ is firing times at which the voltage exceeds $V_{T}$ threshold, $\delta(.)$ is Dirac delta function. Other parameters are constant coefficients, including $\tau_{m}=5ms$ is membrane time scale, $\tau_{w}=100ms$ is adaptation time constant, $R=500M\Omega$ is membrane resistance, $V_{r}=-55mV$ is rest potential, $V_{T}=-51mV$ is voltage threshold, $\Delta_{T}=-2mV$ is sharpness threshold, $a=-0.5ns$ is adaptation voltage coupling, and $b=7pA$ is spike-triggered adaptation increment. At spike time $t=t_{f}$, there will be a sharp rise in the adaptation parameter equal to $b$, as described via the Dirac delta function. The dynamics of the AdEx model are nonlinear and can be chaotic, showing non-smooth deterministic behavior, owing to the threshold-reset mechanism.
Appendix D Comparison of PSD curves in AdEx example
The results presented in FIG. 7 compares the PSD of the membrane potential prediction errors, i.e., $S_{\epsilon_{V}}(f)$. FIG. 9 supplements those results, showing the PSD of the ground truth membrane potential, $S_{V}(f)$, compared to those of the noisy and denoised signals. Note that the difference of the noisy and denoised signals from the ground truth signal is indicated in the PSD curves of FIG. 7 and labeled as “original” and “residual.” These results further confirm the above conclusions, showing the performance of RC in handling different types of noise and emphasizing that the RC is more successful in discarding high-frequency noise.
Figure 9: PSD curves of the membrane potential when considering additive (a) GWN (b) Violet noise (c) Pink noise. “Ground Truth” corresponds to the noise-free original signal; “Noisy” represents the noisy signal fed into the RC framework; and “Denoised” represents the reconstructed signal.